OpenStudy (anonymous):

Find the other complex roots of 4(z^8)+17(z^4)+4=0 if we know 1+i and hence 1-i and 1/(1+i) and 1/(1-i) are solutions.

OpenStudy (anonymous):

\[4z^8+17z^4+4=0\] you wanna find the complex roots as in: \[z^n=r^n(\cos(n\theta)+isin(n\theta))\] ?

OpenStudy (anonymous):

Where is your complex number ? you need to find the roots of ?

OpenStudy (anonymous):

Yeah there should be 8 roots. What do you mean where is my complex number?

OpenStudy (anonymous):

You're given 4 solutions and has to find the 4 remaining. Normally im used to solve for the complex roots when the complex number is given like: a+ib.

OpenStudy (anonymous):

Yes any form is fine

OpenStudy (anonymous):

Im too tired, to help you, im sorry mate : /

OpenStudy (anonymous):

its fine :(

OpenStudy (anonymous):

All right, then, if you know 4 solutions, you can then use polynomial division to extract the last equivalent expression. Try doing that. By this, I mean, for your polynomial \(f(x)\), there is some other polynomial \(g(x)\) such that: \[ f(x)=(x-(1+i))(x-(1-i))(x-(1+i)^{-1})(x-(1-i)^{-1})\;g(x) \]

OpenStudy (anonymous):

Yes/no? Would you like me to expand on this?

OpenStudy (anonymous):

Actually, never mind, I'm blind. What kind of function is your polynomial, check that and you'll have your answer without having to solve anything out.

OpenStudy (anonymous):

so i get \[f(z)=(z^4-3z^3+4.5z^2-3z+1)g(z)\]

OpenStudy (anonymous):

Disregard my previous to last comment. Think of the property of even exponents. A sum of even exponents yields what kind of function?

OpenStudy (anonymous):

Say, we have a function: \[ f(x)=\sum_{n} a_nx^{2n} \]What does that say about: \[ f(-x)? \]

OpenStudy (anonymous):

This is a really pretty problem, and it's a lot easier than it seems; there's very little work involved.

OpenStudy (loser66):

I have a question, please answer me @any mathematicians if I let z^4 = u, then I have 4u^2 +17u+4=0, from then I have u = -0.25 and u = -4 consider u = -0.25 it means z^4 = -0.25 so \(z^2 = \pm0.5\) and continue with the rest wonder whether I am on the right track? appreciate any respond

OpenStudy (anonymous):

does that mean the negative of the roots are also a solutions?

OpenStudy (loser66):

sorry, z^2 =\(\pm0.5i\)

OpenStudy (anonymous):

Yessir. If we have a solution \(f(a)\), then, immediately, \(f(-a)\) is a solution.

OpenStudy (anonymous):

oooh thank you :)

OpenStudy (anonymous):

It's pretty, huh? And, sure thing.

OpenStudy (anonymous):

Here's a little bit to expand: Since \(f(x)\) is an even function, we have: \[ f(x)=f(-x) \]So, for some root \(a\), we have: \[ f(a)=f(-a)=0 \]So, for any root, its negation is immediately a root, itself. Note that this property also applies to odd functions.